# The Effect of Similarity Measures on Accurate Stability Estimates for Local Surrogate Models in Text-based Explainable AI

### v1

@@ 3.3 Success Thresholds
We return to the question of where to stop the adversarial perturbation algorithm to balance fidelity with the original document and a sufficiently different explanation. As what constitutes a sufficiently different explanation is subjective, we instead select a range of values across each of the similarity measures to provide a broad comparison on common choices and to demonstrate the resulting quality of the perturbed document when a larger divergence from the original explanation is required. For our success threshold $\tau$, we choose levels of 30%, 40%, 50%, and 60% for algorithm termination. Thresholds near 70% begin to approach the inherent variation levels of LIME for certain datasets or non-standard sampling rates Burger et al. (2023). For thresholds lower than 30%, the greedy search process is generally too inefficient to find appropriate perturbations for most similarity measures within the maximum perturbation limitations imposed. For measures that require a weighting parameter (RBO) we use the following values 0.5, 0.7, and 0.9. For this formulation values approaching 0 increase the importance of the top features over the remaining features. Values close to 1.0 provide a less concentrated distribution of weight.
@@ 4 Experiment Setup
To provide the raw material for our analysis, we generate batches of 50 adversarial examples using the algorithm in Burger et al. (2023). Each batch is generated with respect to a similarity measure and a success threshold. Nine similarity measures are used: The Jaccard index, Kendall’s Tau Rank Distance, and Spearman’s footrule, each in their standard and weighted implementations and RBO, with weighting parameters 0.5, 0.7, and 0.9. Combined with the thresholds of 30%, 40%, 50%, and 60% this results in 1,800 adversarial examples per dataset. For the datasets we use two of those included in Burger et al. (2023) and their associated pretrained models. The first being the short length (average of 11 words) gender bias Twitter dataset (GB) Dinan et al. (2020) and the second being the moderate length (average of 29 words) symptoms to diagnosis dataset (S2D) (Kaggle). The final and longest length dataset, IMDB movie reviews, proved computationally infeasible due to the excessive time requirements to generate the examples, approximately 125 days of continuous computation on a single A6000. The model to be explained is a DistilBERT Sanh et al. (2019) fine-tuned on each respective dataset. Our interests concern the generalized properties of the similarity measures and not their specific performance associated with a given XAI method and base model and so we choose for computational convenience a single efficient model to explain.
@@ 5 Results and Discussion
All values in the following tables are calculated only for successful attacks. The symbol  -  indicates non-applicability due to no successful attacks on that particular combination of similarity measure, threshold, and dataset.
Attack Success Rate (Table 4): Immediately seen is the inappropriateness of certain similarity measures for use in text-based adversarial XAI attacks. Kendall’s tau is extremely sensitive with nearly 100% attack success across every combination of threshold and dataset. This level of sensitivity renders the measure useless here as the inherent instability of Lime is the cause, little perturbation induced instability is present. This follows for the weighted version as well, though the success rates are not quite as high. Jaccard, Spearman, and Spearmanw also show excessive sensitivity for the higher similarity thresholds of 50% and 60%.
Jaccardw and RBO0.9 exhibit the opposite behavior instead being coarse with few, if any, successful attacks under all but the most lenient threshold. This is not inherently negative, as both measures effectively require the top few features to change, with complete removal from the explanation in the case of Jaccardw and at least a substantial decrease in ranking for RBO0.9. Since much of the weight of an explanation is often associated with the top features, these measures may prove useful as a fast heuristic for a substantial explanation difference.
The attack success rates are generally consistent between the datasets. As the sampling rate for Lime was optimized according to prior work, the inherent instability of the sampling process was kept similar despite the size difference.
Perturbation Quality (Tables 6,7): The quality of the perturbed document follows the expected pattern of more perturbations results in lower quality. We exclude the Kendall measures from our discussion here as they are extremely sensitive and so few perturbations are required to meet the required threshold. Our measures for quality are perplexity (PPL) and similarity as calculated using cosine similarity with the universal sentence encoder (USE).
@@
Minimal Perturbations: There are no strong indications for the choice in similarity measure being a major factor in finding minimal perturbations. We see no average ending similarities appreciably below the success threshold (with the exception of Kendall due to its sensitivity) (Table 5) and there are no notable signs of consistently low perturbation rates (Table 6). However, the search process is very greedy, taking the first viable perturbation that reduces the explanation similarity. An alternative search strategy would likely be superior when focused exclusively on minimal perturbations. Either with a threshold used to specify a minimum reduction in similarity in accepting a perturbation or an entirely new algorithm that can explore the search space more effectively.
@@ 6 Conclusion
We have shown that a poor choice in similarity measure can drastically skew the results of an adversarial attack, either by selecting a measure that is too sensitive and so overstating the weakness of the XAI method, or by choosing a measure too coarse and overstating the method’s resilience. Practitioners and researchers should choose these measures judiciously, keeping in mind how sensitive a measure is to the inherent variation between explanations. Our findings confirm prior work in that measures that use both order and weight produce results with fewer perturbations and subsequently higher textual quality. However, no measure tested proved ideal in selecting for minimal perturbations. It appears necessary that the search process itself will also need adjustment as the current greedy approach is consistent regardless of measure used.
@@
Transferability to other XAI methods: While the fundamentals of the discussion on similarity measures is not unique to Lime, our general conclusions on the fitness of particular similarity measures may not be applicable on different XAI methods. In particular, methods with less inherent instability may see more useful results from measures like Kendall’s Tau where its excessive sensitivity and subsequently exaggerated attack effectiveness may no longer pose a problem.
Computational Expenditure: The computational resources needed to produce the adversarial examples was significant. Existing methods used to generate the attacks were not designed for multi-thousand collections of examples to be generated. Purpose built tools focused on efficiency would allow larger numbers of samples to be reasonably generated which may result in alternate conclusions.
Other similarity measures: The similarity measures used here were designed for the use of comparison of ranked lists. Other measures with this purpose or more generalized measures exist and our conclusions may not be reflective of these untested measures.
@@ Appendix B Bounding Penalized Spearman’s Footrule Maximum Distance
Given a penalty, $p\geq 1$, for disjoint elements (i.e there exists $e\in\mathbf{A}\;s.t\;e\notin\mathbf{B}$), we may have a total possible distance that exceeds the non-adjusted footrule.
@@
As the largest possible individual distance is $|\mathbf{A}|-1$ assume $p\geq\mathbf{|A|}-1$. Then trivially the maximum distance is $p\mathbf{|A|}$.
@@
As $p<\mathbf{|A|}-1$ then there exists some distance $d\;s.t\;p\leq d\leq\mathbf{|A|}-1$. Let this $d$ be located at index $i$ within the pattern $\mathbf{|A|}-1,\mathbf{|A|}-3,\;...$ , where $\mathbf{|A|}-1$ is index $1$. The size of this subset of the above pattern is $\mathbf{|A|}-2i$ and so its sum can be given as $\lfloor\frac{({|\mathbf{A}|-2i})^{2}}{2}\rfloor$.
Then $p_{g}=\lfloor\frac{|\mathbf{A}|^{2}}{2}\rfloor-\lfloor\frac{({\mathbf{{|A|}}-2i)}^{2}}{2}\rfloor$. And since $p\geq d$ then $p_{l}=p(\mathbf{{|A|}}-2i)-\lfloor\frac{({|\mathbf{A}|-2i})^{2}}{2}\rfloor$.
So the maximum distance for $p<\mathbf{|A|}-1$ is given by $p_{g}+p_{l}=\lfloor\frac{|\mathbf{A}|^{2}}{2}\rfloor+p(\mathbf{{|A|}}-2i)-2\lfloor\frac{({|\mathbf{A}|-2i})^{2}}{2}\rfloor$.

### v2

@@ 3.3 Success Thresholds
We return to the question of where to stop the adversarial perturbation algorithm to balance fidelity with the original document and a sufficiently different explanation. As what constitutes a sufficiently different explanation is subjective, we instead select a range of values across each of the similarity measures to provide a broad comparison on common choices and to demonstrate the resulting quality of the perturbed document when a larger divergence from the original explanation is required. For our success threshold $\tau$, we choose levels of 30%, 40%, 50%, and 60% for the completion of the algorithm. Thresholds near 70% begin to approach the inherent variation levels of LIME for certain datasets or nonstandard sampling rates Burger et al. (2023). For thresholds lower than 30%, the greedy search process is generally too inefficient to find appropriate perturbations for most similarity measures within the maximum perturbation limitations imposed. For measures that require a weighting parameter (RBO) we use the following values 0.5, 0.7, and 0.9. For this formulation, values approaching 0 increase the importance of the top features over the remaining features. Values close to 1.0 provide a less concentrated weight distribution.
@@ 4 Experiment Setup
To provide the raw material for our analysis, we generate batches of 50 adversarial examples using the algorithm in Burger et al. (2023). Each batch is generated with respect to a similarity measure and a success threshold. Nine similarity measures are used: The Jaccard index, Kendall’s Tau Rank Distance, and Spearman’s footrule, each in their standard and weighted implementations and RBO, with weighting parameters 0.5, 0.7, and 0.9. Combined with the thresholds of 30%, 40%, 50%, and 60%, this results in 1,800 adversarial examples per data set. For the data sets, we used two of those included in Burger et al. (2023) and their associated pre-trained models. The first is the short length (average of 11 words) gender bias Twitter dataset (GB) Dinan et al. (2020) and the second being the moderate length (average of 29 words) symptoms to diagnosis dataset (S2D) (Kaggle). The final and longest-length dataset, IMDB movie reviews, proved computationally infeasible due to the excessive time requirements to generate the examples, approximately 125 days of continuous computation on a single A6000. The model to be explained is a DistilBERT Sanh et al. (2019) fine-tuned on each respective dataset. Our interests concern the generalized properties of the similarity measures and not their specific performance associated with a given XAI method and base model, and so we choose for computational convenience a single efficient model to explain.
@@ 5 Results and Discussion
All values in the following tables are calculated only for successful attacks. The symbol  -  indicates non-applicability due to the absence of successful attacks on that particular combination of similarity measure, threshold, and data set.
Attack Success Rate (Table 4): Immediately seen is the inappropriateness of certain similarity measures for use in text-based adversarial XAI attacks. Kendall’s tau is extremely sensitive with almost 100% attack success across every combination of threshold and data set. This level of sensitivity renders the measure useless here as the inherent instability of Lime is the cause, and little perturbation-induced instability is present. This follows for the weighted version as well, though the success rates are not quite as high. Jaccard, Spearman, and Spearmanw also show excessive sensitivity for the higher similarity thresholds of 50% and 60%.
Jaccardw and RBO0.9 exhibit the opposite behavior, instead being coarse with few, if any, successful attacks under all but the most lenient threshold. This is not inherently negative, as both measures effectively require the top few characteristics to change, completely removing the explanation in the case of Jaccardw and at least a substantial decrease in the ranking for RBO0.9. Since much of the weight of an explanation is often associated with the top features, these measures may prove useful as a fast heuristic for a substantial explanation difference.
The attack success rates are generally consistent between the datasets. As the sampling rate for Lime was optimized according to previous work, the inherent instability of the sampling process was kept similar despite the size difference.
Perturbation Quality (Tables 6,7): The quality of the perturbed document follows the expected pattern of more perturbations resulting in lower quality. We exclude the Kendall measures from our discussion here as they are extremely sensitive and so few perturbations are required to meet the required threshold. Our quality measures are perplexity (PPL) and similarity calculated using cosine similarity with the universal sentence encoder (USE).
@@
Minimal Perturbations: There are no strong indications for the choice in similarity measure being a major factor in finding minimal perturbations. We see no average ending similarities appreciably below the success threshold (with the exception of Kendall due to its sensitivity) (Table 5) and there are no notable signs of consistently low perturbation rates (Table 6). However, the search process is very greedy, taking the first viable perturbation that reduces the explanation similarity. An alternative search strategy would likely be superior when focused exclusively on minimal perturbations. Either with a threshold used to specify a minimum reduction in similarity in accepting a perturbation or with an entirely new algorithm that can explore the search space more effectively.
@@ 6 Conclusion
We have shown that a poor choice in similarity measure can drastically skew the results of an adversarial attack, either by selecting a measure that is too sensitive and so overstating the weakness of the XAI method, or by choosing a measure too coarse and overstating the method’s resilience. Practitioners and researchers should choose these measures judiciously, keeping in mind how sensitive a measure is to the inherent variation between explanations. Our findings confirm previous work in that measures that use both order and weight produce results with fewer perturbations and subsequently higher textual quality. However, no measure tested proved ideal in selecting for minimal perturbations. It appears necessary that the search process itself will also need adjustment as the current greedy approach is consistent regardless of the measure used.
@@
Transferability to other XAI methods: While the fundamentals of the discussion on similarity measures is not unique to Lime, our general conclusions on the fitness of particular similarity measures may not be applicable on different XAI methods. In particular, methods with less inherent instability may see more useful results from measures like Kendall’s Tau, where its excessive sensitivity and subsequently exaggerated attack effectiveness may no longer pose a problem.
Computational Expenditure: The computational resources needed to produce the adversarial examples were significant. Existing methods used to generate the attacks were not designed for multi-thousand collections of examples to be generated. Purpose-built tools focused on efficiency would allow larger numbers of samples to be reasonably generated, which may result in alternate conclusions.
Other similarity measures: The similarity measures used here were designed for the use of ranked lists. Other measures with this purpose or more generalized measures exist, and our conclusions may not be reflective of these untested measures.
@@ Appendix B Bounding Penalized Spearman’s Footrule Maximum Distance
Given a penalty $p\geq 1$, for disjoint elements (i.e. there exists $e\in\mathbf{A}\;s.t\;e\notin\mathbf{B}$), we may have a total possible distance that exceeds the non-adjusted footrule.
@@
As the largest possible individual distance is $|\mathbf{A}|-1$, assume $p\geq\mathbf{|A|}-1$. Then trivially, the maximum distance is $p\mathbf{|A|}$.
@@
As $p<\mathbf{|A|}-1$, there exists some distance $d\;s.t\;p\leq d\leq\mathbf{|A|}-1$. Let this $d$ be located at the index $i$ within the pattern $\mathbf{|A|}-1,\mathbf{|A|}-3,\;...$ , where $\mathbf{|A|}-1$ is the index $1$. The size of this subset of the above pattern is $\mathbf{|A|}-2i$ and so its sum can be given as $\lfloor\frac{({|\mathbf{A}|-2i})^{2}}{2}\rfloor$.
Then $p_{g}=\lfloor\frac{|\mathbf{A}|^{2}}{2}\rfloor-\lfloor\frac{({\mathbf{{|A|}}-2i)}^{2}}{2}\rfloor$. And since $p\geq d$, then $p_{l}=p(\mathbf{{|A|}}-2i)-\lfloor\frac{({|\mathbf{A}|-2i})^{2}}{2}\rfloor$.
So, the maximum distance for $p<\mathbf{|A|}-1$ is given by $p_{g}+p_{l}=\lfloor\frac{|\mathbf{A}|^{2}}{2}\rfloor+p(\mathbf{{|A|}}-2i)-2\lfloor\frac{({|\mathbf{A}|-2i})^{2}}{2}\rfloor$.